OpenStudy (anonymous):

what ratio is equivalent to 1/7

OpenStudy (anonymous):

2/14

OpenStudy (lgbasallote):

\[\frac 17 \implies \frac{2}{14} \implies \frac{3}{21} \implies \frac{4}{28} \implies \text{etc.}\] if you notice all of them simplifies to 1/7

OpenStudy (zzr0ck3r):

(1(x))/(7(x)) for any x in C

OpenStudy (anonymous):

@sarah43 its only a 14/2

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

@sarah43 my answers are 14/2 3/21 4/11 and 3/14

OpenStudy (anonymous):

3/21 can also be the answer. look @lgbasallote answer

OpenStudy (zzr0ck3r):

there are infinite answers

OpenStudy (lgbasallote):

i already explained how you can determine which one it is ;)

OpenStudy (anonymous):

@igbasallote idkk which one tho im not smart lol

OpenStudy (zzr0ck3r):

any time you have a ratio, if you multiply the top and bottom by the same number then the ratio is perserved. S0 100/700 is answer or pi/(7*pi) or (.0254)/(7*.0254) ...

OpenStudy (zzr0ck3r):